Sol Summary – February 16th

Crew 292 Sol Summary Report 16-02-2024
Sol: Sol 12
Summary Title: Mangalyatri Crew 292 finishes its rotation at MDRS
Author’s name: Annalea Beattie
Mission Status: Nominal
Sol Summary Activity: We finish our simulation today so we are very busy. Clare cooked pancakes and we all were up early(ish). This morning we cleaned and cleaned every corner of the station. We scrubbed everything even under the cupboards in the kitchen. The Science Dome and The Green Hab are spotless. Downstairs the equipment and the area are sparkling. It is incredible really what teamwork can do. Again, we debated where to put biodegradable wipes. Should they be burnt or ? In the end that decision belongs to our engineer who says biodegradable wipes are not paper and are No Burn. That word ‘biodegradable’ is deceptive. Further research and more action needed on this. This morning (after midnight) I also finished our End of Mission Report. We all worked so hard for days to write this and my bad as I missed seeing the word limit on the mission support email. In fact, our actual report with everyone’s research in it is more than twenty pages. So I had to summarise to the short report. We are all happy with the long report however. We decided to turn it into a conference paper as we have really nailed in writing the conceptual breadth and depth of our research here, as well as a real sense of our time together in a simulated Mars environment. What a stunning crew and brilliant work. After lunch we will drive past Skyline Rim to the Valley of the Stars and to the box canyon quarry at White Cliffs, to see at close range authigenic gypsum in the Sommerville.
Look Ahead Plan: Across disciplines we have lots of plans for the Ladakh station to process and a conference paper to submit. We will meet in a few weeks. Mangalyatri Crew 292 will continue to work together. On to Mars Mangalyatri!
Anomalies in work: None
Weather: Early, 51% cloud cover, 30F and high humidity. There was a 5% chance of precipitation. This afternoon the maximum is 51F with 39% humidity.
Crew Physical Status: Nominal
EVA: none
Reports to be filed: Sol Summary, Journalist Report, Green Hab Report, Operations Report, End-of-Mission Report.
Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site